Journalist Report – February 18th

Sol 6 – If the Moon is made of cheese, let Mars be made of brownies…

“He stood looking at the rocket. The ports were open and his crew was streaming out, waving their hands. A crowd of people had gathered, and in and through and among these people the members of the crew were hurrying, talking, laughing, shaking hands. People did little dances. People swarmed.”

– Chapter 6 of The Martian Chronicles by Ray Bradbury

We awoke this morning to a perfectly well-warmed Upper Deck, indicating that the generator had made it through the night and won against the cold! After two days of unusual mornings, we were grateful to go back to our regular schedule, with a full-length workout session devised by our Health and Safety Officer focusing on proprioception, and a smooth daily health check. Jérémy told us tales of his previous missions in Antarctica over breakfast, before getting ready for the weather station EVA. Quentin, Alexandre and Jérémy suited up and headed out with their equipment, took the rovers, and went back to the MegaAres EVA site.

Sitting in the Upper Deck, working as HabCom for today’s EVA crew, I chose today’s quote from The Martian Chronicles in hopes that we would be able to give our EVA crew the proper greeting when they came back to the Hab! With this in mind, Alice prepared a brownie, and Adrien and Corentin returned from the GreenHab to prepare tortillas.

Thanks to Jérémy, Quentin and Alexandre, we now have an up and running weather station. The EVA went smoothly, leaving extra time to run field tests on MegaAres which were mostly successful. We debriefed during lunchtime, before Alice and Jérémy returned to the Science Dome. Alice worked on her project to recreate a 3D geological map using photogrammetry, starting running tests to correctly identify different types of rocks, to then be able to recognize them on the 3D map. “To identify different types of sandstone, you have to look at the grain size,” she explained to me. “I wanted to use the Science Dome microscope to measure them, but had trouble scaling the image I obtained. That will be something to work on next week.”

Meanwhile, in the Hab, Corentin worked on extracting the data from the Polar chest bands we wear at night and during the day. These measure our heartrate and heartrate variability, and also contain an accelerometer to measure our activity. Yesterday, they even allowed us to see how our heart rates were affected by stress related to the power outage. Quentin and Alexandre also had their first EchoFinder session, during which we experienced the same difficulties as yesterday. As I was also in the Hab, I kept switching between touching up photos, trying to capture a timelapse of a Martian sunset and offering them some help.

Tonight, after the com window closes, we plan to celebrate our first week at the MDRS by dancing together and playing some games!

EVA Report – February 18th

Crew 275 EVA 5 Report 18Feb2023
EVA #5

Author: Quentin Royer (Crew Engineer)
Purpose of EVA: Deployment of the meteorological mast and test of Mega-Ares

Start time: 10:03
End time: 12:48

After placing all the equipment needed in the Engineer Airlock, the EVA Crew started their depressurization at 10:03. They departed to the location where the MegaAres antenna was set up during EVA #3, between the MDRS and Marble Ritual, using the Opportunity and Perseverance rovers. The EVA Crew arrived on-site at 10:28.

The EVA Crew unloaded the equipment from the rovers and started to build the weather station and to set up the 3 instruments to be placed on the mast. First, the field mill was set up, but only partially because there was a failure in the battery wire, which will have to be repaired in the RAM. Then, the PurpleAir instrument was set up successfully. Finally, the anemometer was installed, and is fully operating.

The weather station and its instruments were fully deployed at 11:34.

Then, the EVA Crew started performing tests on the MegaAres instrument. Almost all of them were successful, but one of them was a failure, and on repairs were undertaken on-site. They were successful, and the MegaAres instrument started to collect data.

The Crew departed the location at 12:40, returned to the MDRS and unloaded their tools into the Engineer Airlock. The EVA Crew entered the main Airlock at 12:48.

One objective was only partially achieved (Field Mill deployment) and will be finished during another EVA later in the mission.

Destination: Between the Hab and Marble Ritual, near the road

Coordinates (use UTM WGS 84): N4251000 E518500

Participants: EVA leader: Quentin Royer (Executive Officer/Crew Engineer), Jeremy Rabineau (Crew Commander), Alexander Vinas (Crew Astronomer)
HabCom: Marie Delaroche (Crew Journalist)

Road(s) and routes per MDRS Map: Drive Cow Dung Road 0110 to a place just before Marble Ritual

Mode of travel: Driving, even if the destination is less than 1 km away, the crew needed to transport heavy instruments that also take up a lot of space.

GreenHab Report – February 18th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 75.88 °F

Average humidity: 19 % (to be double-checked)

Hours of supplemental light: 4 hours

Daily water usage for crops: 12.68 gallons

Daily water usage for research and/or other purposes: 0 gallons

Water in the blue tank: 121.97 gallons

Time(s) of watering for crops: 11:13 & 16:15

Changes to crops: I relocated the last two tomato plants that were on the upper storage next to the ones I already moved. Indeed, the three I relocated look way better now compared to those which stayed up there.

Narrative: I will shortly double-check the sensor, soon enough the environmental sensors we want to deploy will be ready to use.

Concerning the greenhab, the dirt was a little bit dry on some plants, so I watered them well, but everything looked good. Thus, I maintained and watered crops and plants as my daily operations.

It is a real pleasure to see plants growing every day and to see that the radish I planted pop out of the dirt and come to life.

I decided to relocate the two remaining tomato plants that were on the upper storage to ensure their well-being.

But overall, today was more of a routine than the other days, it means that plants are doing good!

Harvest: N/A

Support/supplies needed: N/A

Operations Report – February 18th

Crew 275 Operations Report 18Feb2023
SOL: 6

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Science Dome: 3 leaks in the walls – same as yesterday
• Propane Generator

Notes on non-nominal systems:
• Concerning toilet, see below in the last section.

Spirit rover used: No
• Hours: 209.0
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Opportunity rover used: Yes
• Hours: 113.1
• Beginning charge: 100%
• Ending charge: 93%
• Currently charging: Yes

Curiosity rover used: No
• Hours: 220.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: Yes
• Hours: 255.5
• Beginning charge: 100%
• Ending charge: 81%
• Currently charging: Yes

General notes on rovers: Rovers were used for today’s EVA. Some snow was removed from the seats of Perseverance.

Summary of Hab operations: The Crew prepared the EVA in the morning and performed routine operations and experiments. The Lower Deck heater was used all day. Temperature in the Lower Deck was at 72°F all day, and the temperature in the Lower Deck was comfortable
• WATER USE: 12.87 gallons
• Water: 383.67 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: No

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS.

Summary of suits and radios: Spacesuits 1, 2 and 3 were used today, nothing to report.

Summary of GreenHab operations: The Crew botanist conducted routine operations, and relocated two tomato crops which were on the upper storage.
• WATER USE: 12.68 gallons
• Heater: On
• Supplemental light: From 10:00 pm to 2:00 am
• Harvest: No

Summary of ScienceDome operations: The Commander and Crew Scientist tested some experimental protocols. The Crew Scientist used a microscope to test a protocol concerning rock samples.
• Dual split: Switched off during the night according to Mission Support

Summary of RAM operations: No operations were conducted in the RAM. We retrieved the 6 AAA batteries and the user manual of the oven.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns, and requests to Mission Support:
• Greenhab radio: the volume of the channel 1 radio was very low. I conducted a check on this radio with the user manual, and the volume is still low. We decided to change radios between channel 1 and 22 in the Greenhab. Indeed, important communication among the Crew are transmitted on channel 1, and transmissions with Mission Support are always managed in the Hab and summarized on channel 1.
• Concerning the toilet, I have listed this system as non-nominal since the beginning of the mission because it was reported as so before. However, we have not experienced and bad smell since the beginning of the mission. Do you still need to change the toilet tank?
• We did not experience any failure for the propane generator, is it still considered as non-nominal?

Sol Summary Report – February 18th

Crew 275 Sol Summary Report 18Feb2023

Sol: 6

Summary Title: If the Moon is made of cheese, let Mars be made of brownies!

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Today, the crew was happy to wake up by the sound of their alarm, rather than by the sound of radio communications with Mission Support. Let’s hope that the temperature will not decrease again and give some very relative rest to the generator (and to Mission Support)! As yesterday’s EVA was postponed to this morning, we already knew the procedures very well and the pre-EVA briefing was relatively short. The three hours of this EVA were used efficiently, as the meteorological mast is now standing proudly right next to Mega-Ares and data acquisition is already ongoing on most of the instruments. In addition, we had time to perform most of the tests on Mega-Ares and launched the acquisition as well, even though some additional tests will be needed during a future EVA (potentially during the deployment of LOAC, that has not been done so far because of the weather conditions). After this successful collaboration between the EVA crew and HabCom, everyone was very happy to share a delicious meal, prepared by the ones who remained at the Hab (we even had brownies!). The other activities continued within the station with absolutely no problem, even though it is always difficult to stick exactly to the hourly schedule meticulously prepared by our Crew Scientist, especially for activities requiring the contribution of several crewmembers.

Look Ahead Plan:

As we are successfully reaching the end of this first week, tomorrow will be a well-deserved rest day for the crew. We will take this opportunity for some team building activities and personal time. This rest time is even more important as week 2 will also be packed with many activities, including the beginning of the photogrammetry EVAs and the end of the deployment of the atmospheric science instruments.

Anomalies in work: None.

Weather: Slight cloud coverage, low -9°C/15°F, high 1°C/34°F.

Crew Physical Status: Optimal.

EVA: None.

Reports to be filed: Sol Summary, EVA Report, Operations Report, Journalist Report, GreenHab Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site